OpenStudy (anonymous):

Opening day attendance at a particular ball field dropped from 45,000 people in 2001 to 31,000 people in 2006. Find the rate of change in opening day attendance.

OpenStudy (anonymous):

The basic principle here is to find the amount of the change and then divide it by the original amount to find the rate of increase or decease. If the result is positive, then it is an increase and if negative, then it is a decrease. (I realize you can tell increase/decrease by the numbers themselves, but am giving you a mathematically sound answer.) 31,000 - 45,000 = -14,000 -14,000/45,000 = 0.68888..... So the percent decrease is approximately 69% If you are looking for a rate as a fraction, it is -14/45.